# PredIN: Towards Open-Set Gesture Recognition via Prediction Inconsistency

### v1

@@ 3.3 Prototype Learning
where $M$ represents the number of known samples. We use the dot product to measure the generalized distance between the samples and prototypes as
$d(\mathbf{z}_{i},\mathbf{p}^{k})=-{\mathbf{z}_{i}}^{T}\cdot{\mathbf{p}^{k}}.$ (4)
@@
Combining Eq. (3) and Eq. (5), the overall loss function $\mathcal{L}_{PL}$ of the PL baseline is expressed as follows:
@@
3.4 Ensemble Class Feature Distribution Inconsistency
In deep learning, class feature distribution is derived from the projection of high-dimensional sample space to low-dimensional feature space with an encoder. Given this, maximizing class feature distribution inconsistency among ensemble members can promote diverse mapping functions, thereby enhancing ensemble diversity. From a global perspective, class feature distribution inconsistency refers to the entirely distinct layout or relative positions of deep features from different classes. Locally, it means that the neighboring classes for each class among ensemble members differ, as illustrated in Fig. 3.
Fig. 3: An illustration of class feature distribution inconsistency between ensemble members. Different colors represent different classes.
Based on PL, the class feature distribution can be approximately characterized by learned prototypes as the actual distribution is not directly accessible and prototypes provide the first-order statistics of the distribution [35]. We compute the distance between each sample feature $\mathbf{z_{i}}$ and prototypes $\mathbf{p}^{j}$ of non-corresponding classes from the same ensemble member:
@@ 3.4 Ensemble Class Feature Distribution Inconsistency
The collection of probabilities represents the class feature distribution within each member, depicted as an $M\times(N-1)$ matrix in Fig. 2. The computation of class feature distribution involves both sample features and prototypes, ensuring that every sample in the feature space contributes to the adjustment in class feature distribution, not just the representatives (prototypes) of feature clusters. To maximize class feature distribution inconsistency, we propose an inconsistency loss which encourages maximizing the inconsistency of two probabilities as follows:
$\mathcal{L}_{incon}=-\frac{1}{M}\sum_{i=1}^{M}\log{\sum_{k=1}^{N}({p_{m_{A},i}^{k}}(1-{p_{m_{B},i}^{k}})+{p_{m_{B},i}^{k}}(1-{p_{m_{A},i}^{k}}))},$ (10)
where $k\neq y_{i}$. Here $m_{A}$ and $m_{B}$ represent two ensemble members, and $p_{m_{A},i}^{k}$ and $p_{m_{B},i}^{k}$ represent the respective probability values in the Eq. (9). From a mathematical perspective, the inconsistency loss is minimized when two probability distributions take opposite extreme values. In terms of class feature distribution, minimizing inconsistency loss $\mathcal{L}_{incon}$ narrows the proximity between two classes within one member while simultaneously increasing the proximity between the corresponding two classes in another member, as shown in Fig. 4.
Fig. 4: An illustration of how inconsistency loss $\mathcal{L}_{incon}$ acts on two members to adjust the class feature distribution. The class pairs of each member are optimized in opposite directions. When the class pairs are pulled within the margin, the optimization of $\mathcal{L}_{incon}$ will halt for one member.
Considering that adjusting the class feature distribution will inevitably pull some features and non-corresponding prototypes close within a member during training, we introduce the positive distance between features and their corresponding prototypes along with a margin $m_{1}>0$ in the distance computation of the Eq. (8) to mitigate this issue:
@@
where $1\leq j\leq N$ and $y_{i}=k\neq j$. The term ${\mathbf{z}_{i}}^{T}\cdot{\mathbf{p}^{k}}$ is not subject to gradient optimization. The redefined distance metric not only considers the relationship with positive pairs ${\mathbf{z}_{i}}$ and ${\mathbf{p}^{k}}$ but also ensures that the inter-class distance does not decrease to within the margin. Specifically, when ${\mathbf{z}_{i}}^{T}\cdot{\mathbf{p}^{j}}>{\mathbf{z}_{i}}^{T}\cdot{\mathbf{p}^{k}}-m_{1}$, the relative positions of corresponding features and prototypes will not be adjusted in one member, as shown in Fig. 4.
3.5 Individual Inter-class Separability
Ensemble learning strategies encourage diversity among members and therefore increase their bias, which may potentially degrade individual performance [15]. While enhancing diversity among ensemble members, we also focus on maintaining individual performance within each member. During maximizing class feature distribution inconsistency, our redefined distance metric ensures a margin for neighboring classes but does not push them away. Since individual ensemble members rely on distance metrics for rejection, establishing a decision boundary with effective inter-class separability is crucial. We therefore introduce the triplet loss [36] based on prototype learning to optimize the inter-class separability. Triplet loss minimizes the distance between an anchor and a positive, both of which belong to the same class, and minimizes the distance between the anchor and a negative of a different class [36]. Neighboring classes naturally form hard negative pairs.
@@ 3.5 Individual Inter-class Separability
where $k=y_{i}$. Class $j$ is the nearest neighbor of class $k$. The loss $\mathcal{L}_{trip}$ pulls the features (anchors) and their corresponding prototypes (positives) closer while pushing them away from the nearest non-corresponding prototypes (negatives), ensuring effective inter-class separability for each individual member.
@@
where $\gamma$ and $\alpha$ are the weights of $\mathcal{L}_{incon}$ and $\mathcal{L}_{trip}$. Here the inconsistency loss and triplet loss are complementary as the former enhances ensemble diversity by maximizing the class feature inconsistency among ensemble members while the latter maintains the individual performance by optimizing the inter-class separability within each member.
3.6 Unknown Rejection
In PredIN, each ensemble member follows previous PL-based approaches and obtains similarity based on the distance between sample features and prototypes. Specifically, similarity between a given feature $\mathbf{z}_{i}$ and the prototype $\mathbf{p}^{k}$ for each member is defined as follows:
@@ 3.6 Unknown Rejection
Fig. 5: Rejection Rules. Through ensemble averaging, unknown samples (right) tend to predict different results and produce lower $S_{max}$, while known samples (left) tend to obtain the same predictions consistent with the label (yellow) and produce higher $S_{max}$.
Besides the above distance metric, we also need to address how to leverage prediction inconsistency brought by ensemble diversity. In ensemble learning, a commonly used method for integrating predictions from different members is averaging [14, 37]. This approach is equally applicable to our task. We combine the outputs of ensemble members:
@@
In summary, we comprehensively consider the distinctions in prediction inconsistency and distance between known and unknown classes, thereby effectively increasing the rejection performance of the unknown. As shown in Fig. 5, in a single member, the maximum similarity score of some unknown samples may be higher than that of known samples, making it difficult to reject based on a certain threshold. However, unknown samples will obtain lower $S_{max}$ through ensemble averaging due to prediction inconsistency, while known samples will be close to the same-class prototypes among ensemble members resulting in higher $S_{max}$. A pre-determined threshold can be applied to $S_{max}$ to reject the unknown. Samples with $S_{max}$ greater than the threshold value will be regarded as known ones and classified into class $k^{*}$.
4 Experiments and results
4.1 Datasets
We apply four public sEMG benchmark datasets [38, 18, 39, 40] to validate the proposed approach, as shown in Table 1. During preprocessing, raw sEMG signals are segmented via a sliding window of length 200 ms with steps of 50 ms, and then standardized channel-wise. As recommended by BioPatRec [38], we remove transient periods of the contraction using a contraction time percentage of $0.7$ for the BioPat DB2 dataset. According to the setting of closed-set gesture recognition based on sEMG [2, 3, 4], the training and testing set are divided based on trials as mentioned in Table 1. Following the protocol of open-set image recognition [13], we randomly select $10$ known classes from BioPat DB2 and $15$ known classes from Ninapro DB2, Ninapro DB4 and Ninapro DB7, while treating the remaining classes as unknown.
Table 1: Characteristics and setup of four public sEMG datasets.
Dataset BioPat DB2 Ninapro DB2 Ninapro DB4 Ninapro DB7
@@ 4.1 Datasets
4.2 Evaluation Metrics
@@
4.3 Experimental Settings
To comprehensively extract sEMG features, we employ two types of encoders, the Crossformer [41] and a CNN-LSTM hybrid network, as the backbone. The Crossformer is a popular Transformer-based model for time series forecasting while also showing superiority in sEMG classification tasks since it effectively captures the cross-time and cross-channel dependency and extracts multi-scale time information. We set the segment length of Crossformer as $32$, window size as $2$, and layers as $5$. In addition, we design a hybrid network based on CNN and LSTM, which combines a ResNet variant [42], an LSTM [43] and an SKAttention module [44].
During experiments, we use the SGD optimizer with an initial learning rate of $0.01$. The learning rate decreases by a factor of $0.1$ at $60$ and $80$ epochs. The batch size is set to $256$ and the training epoch is set to $100$. The hyperparameters $\beta$, $\gamma$ and $\alpha$ in Eq. (7) and Eq. (13) are all empirically set to $1.0$, while the feature dimension of embedding space is set to $128$. Two margins in Eq. (11) and Eq. (12) are set to $0.5$ and $1.0$ respectively. Prototypes are randomly initialized by the standard normal distribution. All experimental results are averaged among five randomized splits of datasets by classes, which means each split uses different known classes to train. The PredIN is implemented by using Pytorch 2.3.0 and executed on an NVIDIA GeForce RTX 4090 GPU.
Table 2: Performance comparison with SOTAs in terms of AUC (%) and OSCR (%) on four public datasets. Results are averaged among five randomized splits. Best performances are highlighted in bold.
@@ 4.3 Experimental Settings
 AUC OSCR  AUC OSCR  AUC OSCR  AUC OSCR
Softmax 65.5 60.5 72.2 65.6 73.1 61.7 73.5 66.8
ARPL (TPAMI’22) 68.2 61.1 76.3 67.9 79.6 64.5 79.1 70.1
SLCPL (CVIU’23) 70.1 63.7 76.5 68.4 76.4 63.3 76.6 68.4
DIAS (ECCV’22) 70.2 65.8 77.8 69.7 80.2 65.7 79.8 71.0
CPN-MGR (IEEE Sens. J’22) 69.7 63.3 76.4 68.5 76.6 63.2 76.4 68.4
PredIN (Crossformer) 74.9 69.8 80.4 72.5 80.8 68.2 81.0 73.1
OpenMax (CVPR’16) 68.6 64.7 69.2 58.4 74.5 61.2 71.1 62.2
ARPL (TPAMI’22) 73.4 70.0 72.2 59.1 79.8 63.0 78.6 64.6
MGPL (Inf. Sci’23) 70.6 67.3 62.3 44.7 76.7 58.3 64.9 51.7
MEDAF (AAAI’24) 74.7 72.2 72.9 63.2 80.2 64.9 79.9 70.5
CPN-MGR (IEEE Sens. J’22) 73.0 69.3 71.9 55.4 79.3 58.0 78.2 63.6
PredIN (Hybrid model) 75.7 72.7 76.7 64.3 82.2 67.6 82.1 71.6
Table 3: Performance comparison with SOTAs in terms of closed-set ACC (%) on four public datasets. Results are averaged among five randomized splits. Best performances are highlighted in bold.
Methods  BioPat DB2 Ninapro DB2 Ninapro DB4 Ninapro DB7
Softmax 86.0 82.6 75.3 83.0
ARPL (TPAMI’22) 85.7 82.7 74.8 82.9
SLCPL (CVIU’23) 85.8 82.5 75.5 82.8
DIAS (ECCV’22) 88.1 83.5 75.9 83.5
CPN-MGR (IEEE Sens. J’22) 85.9 82.6 75.0 83.1
PredIN (Crossformer) 89.2 85.2 78.1 85.2
OpenMax (CVPR’16) 87.8 74.7 73.0 78.2
ARPL (TPAMI’22) 91.9 74.6 73.1 78.5
MGPL (Inf. Sci’23) 91.5 65.4 70.1 68.9
MEDAF (AAAI’24) 93.0 77.7 74.4 82.4
CPN-MGR (IEEE Sens. J’22) 92.0 69.5 67.7 75.3
PredIN (Hybrid model) 93.4 77.6 76.9 82.2
4.4 Comparison with the State-of-the-arts
We compare our method against other state-of-the-art open-set image and gesture recognition approaches. Softmax and OpenMax are methods based on the prediction probability. OpenMax [21] uses extreme value theory (EVT) to calibrate the prediction probability. ARPL [13] and SLCPL [23] are methods based on prototype learning and design various distance loss functions to reduce open space risk. DIAS [26] considers different difficulty levels and introduces an image generator and a feature generator to produce hard fake instances. MGPL [25] is also the method based on PL but applies the VAE framework to optimize generative constraints. MEDAF [28] is the method based on ensemble learning, which encourages multiple experts to learn diverse representation with an attention diversity regularization. CPN-MPR [7] focuses on open-set sEMG-based gesture recognition and introduces the PL to reject the unknown. To ensure the fairness of the comparison, all methods employ the same backbone.
The results in Table 2 and Table 3 highlight the performance of our proposed approach for the open-set sEMG-based gesture recognition task, even though across different backbone architectures. Specifically, considering the rejection performance, our method achieves the best AUC scores of 75.7%, 76.7%, 82.2% and 82.1% across four datasets. Moreover, ensemble learning brings benefits to closed-set classification tasks. Our method achieves significant improvements in closed-set accuracy compared to these SOTA methods. Furthermore, when considering both empirical classification risk and open space risk, our approach also surpasses the above SOTA methods, consistently achieving the highest OSCR scores on four datasets. These results confirm that prediction inconsistency reveals the distinctions between known and unknown classes effectively. In conclusion, our approach shows the superiority in both closed-set classification and unknown rejection.
Table 4: Ablations of each module in terms of AUC (%) and OSCR (%) on four public datasets. Best performances are highlighted in bold.
Methods BioPat DB2 Ninapro DB2 Ninapro DB4 Ninapro DB7
 AUC OSCR  AUC OSCR  AUC OSCR  AUC OSCR
PL baseline 71.5 66.5 73.3 59.1 79.1 61.7 77.7 65.4
Deep Ensemble 73.2 68.9 76.0 64.0 80.9 65.8 80.2 70.2
Deep Ensemble (w/ $\mathcal{L}_{trip}$) 73.8 70.7 76.4 64.2 81.8 67.4 80.8 71.0
PredIN (w/o $\mathcal{L}_{trip}$) 72.4 68.3 75.6 61.9 81.6 66.4 81.1 70.6
PredIN 75.7 72.7 76.7 64.3 82.2 67.6 82.1 71.6
4.5 Ablation Study
Module Ablation. As presented in Table 4, each component within our method has been systematically integrated into the PL baseline to verify its necessity. We first introduce the Deep Ensemble framework, which combines two identical networks with different random initializations. As shown in the second raw, the introduction of ensemble learning yields notable improvements in AUC and OSCR scores compared to the PL baseline, demonstrating the effectiveness of using prediction inconsistency for unknown rejection. Secondly, we further verify the effectiveness of our proposed two losses. In PredIN, the roles of inconsistency loss and triplet loss are complementary: the loss $\mathcal{L}_{trip}$ acts to optimize the inter-class separability with each ensemble member while the inconsistency loss $\mathcal{L}_{incon}$ maximizes the class feature distribution inconsistency among members. To further explain, we apply the loss $\mathcal{L}_{trip}$ to the Deep Ensemble model alone and then combine these two losses into the model training. The standalone application of $\mathcal{L}_{trip}$ enhances the individual rejection capability, thereby consistently improving overall ensemble performance compared to the Deep Ensemble model. When combining these two losses, the PredIN magnifies the prediction inconsistency by enhancing the ensemble diversity while maintaining the individual performance, which provides further improvements and clearly demonstrates the effectiveness of maximizing the class feature distribution inconsistency. In addition, we remove the loss $\mathcal{L}_{trip}$ from the PredIN to verify the importance of maintaining individual performance from another perspective. A clear decrease occurs on BioPat DB2 and Ninapro DB2 after the removal. Finally, incorporating all the above components improves AUC by an average of +3.8% compared to the baseline PL model, which demonstrates that each component contributes to the overall improvement on unknown rejection.
Hyperparameters Ablation. We evaluate the effect of two sets of hyperparameters: the trade-off weights $\beta$ and $\gamma$ in Eq. (7) and Eq. (13), and two margins $m_{1}$ and $m_{2}$ in Eq. (11) and Eq. (12). These experiments are conducted using the bioDB2 dataset. Fig. LABEL:fig6:a shows the AUC for different $\beta$ and $\gamma$ values. The weight $\beta$ influences the compactness of feature space in prototype learning. According to the results, larger values yield better results, but excessively large values may cause optimization issues and lead to a performance decrease. The weight $\gamma$ represents the degree of adjustment to the class feature distribution. Larger values similarly affect the model’s classification. Setting $\beta$ and $\gamma$ to $1.0$ is optimal. Two margins influence the degree of class feature distribution inconsistency and inter-class separability. Greater values of $m_{1}$ and $m_{2}$ lead to reduced inconsistency but increased enhancement for inter-class separability. A balanced result is that the optimal values for $m_{1}$ and $m_{2}$ are set to $0.5$ and $1.0$ in our experiments, respectively.
5 Further Analysis and Discussion
5.1 Image Domain Verification
Table 5: Performance comparison (%) with SOTAs in terms of AUC on four public image datasets. Results are averaged among five randomized splits. Best performances are highlighted in bold. * indicates the reproduced result to unify the split information.
Methods MNIST SVHN CIFAR10 TinyIN
Softmax 97.8 88.6 67.7 57.7
OpenMax (CVPR’16) 98.1 89.4 69.5 57.6
CROSR (CVPR’19) 99.1 89.9 88.3 58.9
PROSER (CVPR’21) - 94.3 89.1 69.3
CPN (TPAMI’22) 99.0 92.6 82.8 63.9
ARPL (TPAMI’22) 99.6 96.3 90.1 76.2
ODL (TPAMI’22) 99.6 95.4 88.5 74.6
SLCPL (CVIU’23) 99.4 95.2 86.1 74.9
MGPL (Inf.Sci’23) - 95.7 84.0 73.0
m-OvR (Pattern Recognit.’24) - 95.7 89.5 75.3
DIAS* (ECCV’22) 99.5 94.7 90.3 76.8
PredIN 99.6 97.2 90.5 77.2
To further verify the performance of our proposed approach, we make a comparative experiment on four public image datasets widely used for OSR performance evaluation, including MNIST, SVHN [45], CIFAR10 [46] and TinyImageNet [47]. We compare our proposed method, PredIN, to the SOTA open-set recognition methods including Softmax, OpenMax [21], CROSR [48], PROSER [22], CPN [12], ARPL [13], ODL [49], SLCPL [23], MGPL [25], m-OvR [11] and DIAS [26]. For a fair comparison, we use the same backbone VGG [50] as these methods. In terms of optimization, we use the SGD optimizer with a momentum value of $0.9$ and set the initial learning rate to $0.01$ which drops to $0.1$ at every 30 epochs. The parameters $\beta$, $\gamma$, $\alpha$, $m_{1}$, $m_{2}$ and feature dimension are set to $1.0$, $1.0$, $1.0$, $0.5$, $0.5$ and $128$, respectively. All results of SOTA methods are taken from the references except DIAS [26]. As DIAS [26] applies different dataset split ways, we reproduce their method using the recommended hyperparameters to unify the split information. AUC performances are shown in Table 5. Our approach achieves comparable performance to SOTA methods. This clearly demonstrates the general applicability of our approach across sEMG and image domains.
(a) Deep Ensemble model

### v2

@@ III-C Prototype Learning
where $M$ represents the number of known samples. We use the dot product to measure the generalized distance between the samples and prototypes.
@@
Combining (3) and (4), the overall loss function $\mathcal{L}_{PL}$ of the PL baseline is expressed as follows:
@@
III-D Dual-perspective Class Feature Distribution Inconsistency
In deep learning, class feature distribution is derived from the projection of high-dimensional sample space to low-dimensional feature space with an encoder. Given this, maximizing class feature distribution inconsistency among different perspectives can promote diverse mapping functions, thereby enhancing their differences. Globally, class feature distribution inconsistency refers to the entirely distinct layout or relative positions of deep features from different classes. Locally, it means that the neighboring classes for each class among branches differ, as illustrated in Figure 4.
Figure 4: An illustration of class feature distribution inconsistency between dual perspectives. Different colors represent different classes.
Based on PL, the class feature distribution can be approximately characterized by learned prototypes since the actual distribution is not directly accessible and prototypes provide the first-order statistics of the distribution [32]. We compute the distance between each sample feature $\mathbf{z_{i}}$ and prototypes $\mathbf{p}^{j}$ of non-corresponding classes from the same branch:
@@ III-D Dual-perspective Class Feature Distribution Inconsistency
The collection of probabilities represents the class feature distribution within each branch, depicted as an $M\times(N-1)$ matrix in Figure 3. The computation of class feature distribution involves both sample features and prototypes, ensuring that every sample in the feature space contributes to the adjustment in class feature distribution, not just the representatives (prototypes) of feature clusters. To maximize class feature distribution inconsistency, we propose an inconsistency loss which encourages maximizing the inconsistency of two probabilities as follows:
$\mathcal{L}_{incon}=-\frac{1}{M}\sum_{i=1}^{M}\log$ $\sum_{k=1}^{N}({p_{Br_{A},i}^{k}}(1-{p_{Br_{B},i}^{k}})$ (9)
$+{p_{Br_{B},i}^{k}}(1-{p_{Br_{A},i}^{k}})),$
where $k\neq y_{i}$. Here $Br_{A}$ and $Br_{B}$ represent two branches, and $p_{Br_{A},i}^{k}$ and $p_{Br_{B},i}^{k}$ represent the respective probability values in the (8). From a mathematical perspective, the inconsistency loss is minimized when two probability distributions take opposite extreme values. In terms of class feature distribution, minimizing inconsistency loss $\mathcal{L}_{incon}$ narrows the proximity between two classes within one branch while simultaneously increasing the proximity between the corresponding two classes in another branch, as shown in Figure 5.
Figure 5: An illustration of how inconsistency loss $\mathcal{L}_{incon}$ acts on two branches to adjust the class feature distribution. The class pairs of each branch are optimized in opposite directions. When the class pairs are pulled within the margin, the optimization of $\mathcal{L}_{incon}$ will halt for one branch.
Considering that adjusting the class feature distribution will inevitably pull some features and non-corresponding prototypes close within a branch during training, we introduce the positive distance between features and their corresponding prototypes along with a margin $m_{1}>0$ in the distance computation of the (7) to mitigate this issue:
@@
where $1\leq j\leq N$ and $y_{i}=k\neq j$. The term ${\mathbf{z}_{i}}^{T}\cdot{\mathbf{p}^{k}}$ is not subject to gradient optimization. The redefined distance metric not only considers the relationship with positive pairs ${\mathbf{z}_{i}}$ and ${\mathbf{p}^{k}}$ but also ensures that the inter-class distance does not decrease to within the margin. Specifically, when ${\mathbf{z}_{i}}^{T}\cdot{\mathbf{p}^{j}}>{\mathbf{z}_{i}}^{T}\cdot{\mathbf{p}^{k}}-m_{1}$, the relative positions of corresponding features and prototypes will not be adjusted in one branch, as shown in Figure 5.
III-E Individual Inter-class Separability
Dual-perspective inconsistency learning encourages differences among perspectives and therefore increase their bias, which may potentially degrade individual performance. While enhancing differences among the dual perspectives, we also focus on maintaining individual performance within each branch. Since individual branches rely on distance metrics for rejection, establishing a decision boundary with effective inter-class separability is crucial. We therefore introduce the triplet loss [33] based on prototype learning to optimize the inter-class separability. Triplet loss minimizes the distance between an anchor and a positive, both of which belong to the same class, and minimizes the distance between the anchor and a negative of a different class [33]. Neighboring classes naturally form hard negative pairs.
@@ III-E Individual Inter-class Separability
where $k=y_{i}$. Class $j$ is the nearest neighbor of class $k$. The loss $\mathcal{L}_{trip}$ pulls the features (anchors) and their corresponding prototypes (positives) closer while pushing them away from the nearest non-corresponding prototypes (negatives), ensuring effective inter-class separability for each individual branch.
@@
where $\gamma$ and $\alpha$ are the weights of $\mathcal{L}_{incon}$ and $\mathcal{L}_{trip}$. Here the inconsistency loss and triplet loss are complementary as the former enhances differences within perspectives by maximizing the class feature distribution inconsistency between branches while the latter maintains the individual performance by optimizing the inter-class separability within each branch.
Figure 6: Rejection rules. Through averaging, unknown samples (right) tend to predict different results and produce lower $S_{max}$, while known samples (left) tend to obtain the same predictions consistent with the label (yellow) and produce higher $S_{max}$.
III-F Unknown Rejection
In PredIN, each branch follows previous PL-based approaches and obtains similarity based on the distance between sample features and prototypes. Specifically, similarity between a given feature $\mathbf{z}_{i}$ and the prototype $\mathbf{p}^{k}$ for each branch is defined as follows:
@@ III-F Unknown Rejection
Besides the above distance metric, we also need to address how to leverage prediction inconsistency brought by different perspectives. A commonly used method for integrating predictions from different branches is averaging, which is equally applicable to our task. We combine the outputs of two branches:
@@
TABLE I: Characteristics and setup of four public sEMG datasets. BioDB2, DB2, DB4, and DB7 are abbreviations for BioPat DB2, Ninapro DB2, DB4, and DB7, respectively.
Dataset BioDB2 DB2 DB4 DB7
@@
In summary, we comprehensively consider the distinctions in prediction inconsistency and distance between known and unknown classes, thereby effectively increasing the rejection performance of the unknown. As shown in Figure 6, in a single branch, the maximum similarity score of some unknown samples may be higher than that of known samples, making it difficult to reject based on a certain threshold. However, unknown samples will obtain lower $S_{max}$ through averaging due to prediction inconsistency, while known samples will be close to the same-class prototypes among different perspectives resulting in higher $S_{max}$. A pre-determined threshold can be applied to $S_{max}$ to reject the unknown. Samples with $S_{max}$ greater than the threshold value will be regarded as known ones and classified into class $k^{*}$. Building on the majority of previous OSR studies, a threshold that ensures 95% of the data is recognized as known is commonly used.
IV Experiments and results
IV-A Datasets
TABLE II: Performance comparison with SOTAs in terms of AUC (%), OSCR (%) and ACC (%) on four public datasets. Results are averaged among five randomized splits. Best performances are highlighted in bold.
Methods BioPat DB2 Ninapro DB2 Ninapro DB4 Ninapro DB7
 AUC OSCR  ACC  AUC OSCR  ACC  AUC OSCR  ACC  AUC OSCR  ACC
Softmax 65.5 60.5 86.0 72.2 65.6 82.6 73.1 61.7 75.3 73.5 66.8 83.0
PROSER (CVPR’21) 67.9 63.5 87.8 72.4 63.9 80.0 74.6 63.1 76.1 73.5 66.7 82.9
CAC (WACV’21) 68.6 63.9 87.7 72.9 65.8 82.5 75.5 63.9 76.4 72.6 66.3 83.7
CPN-MGR (IEEE Sens. J’22) 69.7 63.3 85.9 76.4 68.5 82.6 76.6 63.2 75.0 76.4 68.4 83.1
ARPL (TPAMI’22) 68.2 61.1 85.7 76.3 67.9 82.7 79.6 64.5 74.8 79.1 70.1 82.9
SLCPL (CVIU’23) 70.1 63.7 85.8 76.5 68.4 82.5 76.4 63.3 75.5 76.6 68.4 82.8
DIAS (ECCV’22) 70.2 65.8 88.1 77.8 69.7 83.5 80.2 65.7 75.9 79.8 71.0 83.5
PredIN (Crossformer) 74.9 69.8 89.2 80.4 72.5 85.2 80.8 68.2 78.1 81.0 73.1 85.2
OpenMax (CVPR’16) 68.6 64.7 87.8 69.2 58.4 74.7 74.5 61.2 73.0 71.1 62.2 78.2
PROSER (CVPR’21) 75.0 72.6 93.5 69.6 56.1 71.4 76.2 59.3 69.9 73.4 63.6 78.7
CAC (WACV’21) 72.2 69.9 93.1 70.4 58.8 74.7 76.8 63.1 73.8 73.2 64.0 79.5
CSSR (TPAMI’22) 70.0 65.1 92.4 70.3 56.9 72.7 77.7 57.5 73.5 73.3 56.7 78.0
CPN-MGR (IEEE Sens. J’22) 73.0 69.3 92.0 71.9 55.4 69.5 79.3 58.0 67.7 78.2 63.6 75.3
ARPL (TPAMI’22) 73.4 70.0 91.9 72.2 59.1 74.7 79.8 63.0 73.1 78.6 64.6 78.2
MGPL (Inf. Sci’23) 70.6 67.3 91.5 62.3 44.7 65.4 76.7 58.3 70.1 64.9 51.7 68.9
SLCPL (CVIU’23) 72.1 69.5 93.2 71.5 57.3 75.8 77.9 61.8 72.3 77.1 64.1 76.8
MEDAF (AAAI’24) 74.7 72.2 93.0 72.9 63.2 77.7 80.2 64.9 74.4 79.9 70.5 82.4
PredIN (Hybrid model) 75.7 72.7 93.4 76.7 64.3 77.6 82.2 67.6 76.9 82.1 71.6 82.2
We apply four public sEMG benchmark datasets [34, 18, 35, 36] to validate the proposed approach, as shown in Table I. During preprocessing, raw sEMG signals are segmented via a sliding window of length 200 ms with steps of 50 ms, and then standardized channel-wise. As recommended by BioPatRec [34], we remove transient periods of the contraction using a contraction time percentage of $0.7$ for the BioPat DB2 dataset. According to the setting of closed-set gesture recognition based on sEMG [19, 37], the training and testing set are divided based on trials as mentioned in Table I. Following the protocol of open-set image recognition [14], we randomly select $10$ known classes from BioPat DB2 and $15$ known classes from Ninapro DB2, Ninapro DB4 and Ninapro DB7, while treating the remaining classes as unknown.
IV-B Evaluation Metrics
@@
IV-C Experimental Settings
To comprehensively extract sEMG features, we employ two types of encoders, the Crossformer [38] and a CNN-LSTM hybrid network, as the backbone. The Crossformer is a popular Transformer-based model for time series forecasting while also showing superiority in sEMG classification tasks since it effectively captures the cross-time and cross-channel dependency and extracts multi-scale time information. We set the segment length of Crossformer as $32$, window size as $2$, and layers as $5$. In addition, we design a hybrid network based on CNN and LSTM, which combines a ResNet variant, an LSTM and an SKAttention module.
During experiments, we use the SGD optimizer with an initial learning rate of $0.01$. The learning rate decreases by a factor of $0.1$ at $60$ and $80$ epochs. The batch size is set to $256$ and the training epoch is set to $100$. The hyperparameters $\beta$, $\gamma$ and $\alpha$ in (6) and (12) are all empirically set to $1.0$, while the feature dimension of embedding space is set to $128$. Two margins in (10) and (11) are set to $0.5$ and $1.0$ respectively. Prototypes are randomly initialized by the standard normal distribution. All experimental results are averaged among five randomized splits of datasets by classes, which means each split uses different known classes to train. The PredIN is implemented by using Pytorch 2.3.0 and executed on an NVIDIA GeForce RTX 4090 GPU.
IV-D Comparison with the State-of-the-arts
We compare our method against other state-of-the-art open-set image and gesture recognition approaches. Softmax and OpenMax are methods based on the prediction probability. OpenMax [22] uses extreme value theory (EVT) to calibrate the prediction probability. PROSER [23] augmented the closed-sed classifier with an extra classifier placeholder and mimics new data with manifold mixup. CAC [39] proposes a new distanced-based loss function to encourage clustering among known classes. CSSR [11] replaces prototype points with manifolds represented by class-specific auto-encoders(AE). ARPL [14] and SLCPL [26] are methods based on prototype learning and design various distance loss functions to reduce open space risk. DIAS [30] considers different difficulty levels and introduces an image generator and a feature generator to produce hard fake instances. MGPL [28] is also the method based on PL but applies the VAE framework to optimize generative constraints. MEDAF [29] learns diverse representation with an attention diversity regularization. CPN-MPR [9] focuses on open-set sEMG-based gesture recognition and introduces the PL to reject the unknown. To ensure the fairness of the comparison, all methods employ the same backbone.
The results in Table II highlight the performance of our proposed approach for the open-set sEMG-based gesture recognition task, even though across different backbone architectures. Specifically, considering the rejection performance, our method achieves the best AUC scores of 75.7%, 76.7%, 82.2% and 82.1% across four datasets. Moreover, dual-perspective inconsistency learning brings benefits to closed-set classification tasks. Our method achieves significant improvements in closed-set accuracy compared to these SOTA methods. Furthermore, when considering both empirical classification risk and open space risk, our approach also surpasses the above SOTA methods, consistently achieving the highest OSCR scores on four datasets. These results confirm that prediction inconsistency reveals the distinctions between known and unknown classes effectively. In conclusion, our approach shows the superiority in both closed-set classification and unknown rejection.
TABLE III: Ablations of each module in terms of AUC (%), OSCR (%) and ACC (%) on four public datasets. Best performances are highlighted in bold.
@@ IV-D Comparison with the State-of-the-arts
 AUC OSCR  ACC  AUC OSCR  ACC  AUC OSCR  ACC  AUC OSCR  ACC
PL baseline 71.5 66.5 89.1 73.3 59.1 73.2 79.1 61.7 71.9 77.7 65.4 77.7
Dual-perspective 73.2 68.9 90.4 76.0 64.0 77.5 80.9 65.8 75.5 80.2 70.2 81.8
Dual-perspective (w/ $\mathcal{L}_{trip}$) 73.8 70.7 92.3 76.4 64.2 77.4 81.8 67.4 76.9 80.8 71.0 82.2
PredIN (w/o $\mathcal{L}_{trip}$) 72.4 68.3 90.3 75.6 61.9 75.6 81.6 66.4 75.6 81.1 70.6 81.8
PredIN 75.7 72.7 93.4 76.7 64.3 77.6 82.2 67.6 76.9 82.1 71.6 82.2
Figure 7: Evaluation on hyperparamers impact. (a) The AUC performance under different weights $\beta$ and $\gamma$. (b) The AUC performance under different margins $m_{1}$ and $m_{2}$.
(a) Dual-perspective